# Driving of an open quantum system at finite temperature across first- and second-order quantum phase transitions

### v2

@@ 4.2 Second-order QPT
The peak of some fidelity curves, see item (e), has been subject to the same analysis as in the first-order QPT case, the result being summarized in Fig. 8. We again observe the linear increase of the optimal temperature $T_{\rm opt}$ and an roughly an algebraic decrease of ${\rm max}\,\mathcal{F}_{2}$ with the size $N$ of the qubit system.
In the case of the second-order QPT path, special attention needs to be payed to the choice of the system-environment coupling operator $\hat{Q}$. As pointed out in item (f), this choice makes a large difference in the values of $\mathcal{F}_{2}$ obtained for the strong coupling. This effect is due to the difference between both choices in terms of the parity conservation. We know that for ${q=0}$ the driving from the positive-parity initial state along the ${\chi=0}$ path can excite the system only to states with positive parity. However, for ${q\neq 0}$ the parity may be violated by system-environment interaction. This happens for ${\hat{Q}=\hat{J}_{x}}$, so during the drive the environment induces strong transitions to the first excited state with negative parity. This population is then counted in the summed final fidelity $\mathcal{F}_{2}$ in Eq. (3.3), so for slow-enough driving (when transitions to higher excited states can be neglected) this fidelity may be relatively high.
In contrast, for ${\hat{Q}=\hat{J}_{z}}$ the parity is conserved even by the system-environment interaction, so the strongest transitions during the drive lead to the second excited state with positive parity, which does not contribute to $\mathcal{F}_{2}$. This explains why slow drivings in the strong-coupling columns of Fig. 7 yield higher fidelity for ${\hat{Q}=\hat{J}_{x}}$ than for ${\hat{Q}=\hat{J}_{z}}$. In the latter case all $\mathcal{F}_{2}$ curves converge to a narrow band, washing out the difference between slow and fast driving. It seems that for ${\hat{Q}=\hat{J}_{z}}$ the environment-induced transitions and the transitions of the Landau-Zener type have similar effects which complement each other as the driving time $t_{\rm F}$ varies.
4.3 Comparison with the Lindblad method
In this section we compare the above HEOM results with simulations using the Lindblad method. The use of the Lindblad formalism for time-dependent Hamiltonians was discussed, e.g., in Refs. Alba12 ; Yama17 ; Dann18 . In our calculations we apply the Lindblad formula
$\frac{d}{dt}\hat{\rho}_{\text{S}}(t)=-i[\hat{H}_{\text{S}}(t),\hat{\rho}_{\text{S}}(t)]+\mathcal{D}_{t}\hat{\rho}_{\text{S}}(t)$ (29)
with the dissipator adapted from Ref. Yama17 ,
[…]
Here, the variable ${\epsilon=E_{m}(\mathbb{\Lambda}(t))-E_{n}(\mathbb{\Lambda}(t))}$ denotes all possible differences between instantaneous system energies (we drop the dependence on $\mathbb{\Lambda}$) and the operator
$\hat{S}(\epsilon)=\sum\limits_{\begin{subarray}{c}m,n\\ E_{m}-E_{n}=\epsilon\end{subarray}}\langle E_{n}|\hat{Q}|E_{m}\rangle|E_{n}\rangle\langle E_{m}|.$ (31)
The function $\Gamma(\epsilon)$ denotes the one-sided Fourier transform of the bath correlation function, which for Drude-Lorentz distribution (11) takes the form
$\Gamma(\epsilon)=\pi\frac{c_{0}}{\gamma-i\epsilon}+\pi\sum_{k=1}\frac{c_{k}}{2\pi kT-i\epsilon},$ (32)
with coefficients $c_{0},c_{1},\ldots$ introduced in Eq. (26).
@@ 4.3 Comparison with the Lindblad method
Figure 9: Fidelity for driving A across the first-order QPT for the ${N=10}$ qubit system obtained by the HEOM method (full curves) and by the Lindblad method according to Ref. Yama17 (dashed curves). The top row corresponds to driving time ${t_{\rm F}=10^{3.6}}$, the bottom row to ${t_{\rm F}=10^{2.8}}$. The left and right columns correspond to weak and strong system-environment couplings, respectively.
@@
In the comparison of the Lindblad calculations with the HEOM results, we consider only the path across the first-order QPT with the driving protocol A (see Sec. 4.1). If the dynamics for ${N=10}$ qubits is modeled using the two-state approximation, the adiabatic timescale is of the order ${t=2|d\Delta_{10}/ds|/\pi\Delta_{10}^{2}\sim 10^{3}}$, where all terms are evaluated at the crossing Dick13 . The method of Ref. Alba12 ; Yama17 is limited to nearly-adiabatic drivings, so only the drivings with ${t_{\rm F}\gg 10^{3}}$ can be expected to agree with the Lindblad calculations. In Fig. 9 we show the fidelity $\mathcal{F}_{1}$ from Eq. (24) obtained by both the HEOM and Lindblad methods for driving times $t_{\rm F}=10^{2.8}$ and $10^{3.6}$. We see, in accord with the above expectation, that for the larger $t_{\rm F}$ the discrepancy between both methods affects only the drivings at lower temperatures ($T\lesssim\gamma/2\pi$, see Ref. Tani20 ), which shows up only for the strong system-environment coupling. In contrast, for the shorter $t_{\rm F}$ the HEOM and Lindblad results differ significantly even for the weak coupling, both calculations coinciding only when the fidelity approaches the high-$T$ limit (27).
@@
Therefore we can conclude that non-Markovian effects, properly captured in our HEOM calculations, play in general an important role, which is not reproduced by the simpler Lindblad calculations.
@@ 5 Conclusion
However, thermalization and dissipation processes can strongly modify results of the driving procedure, in some cases being able to considerably improve its fidelity. One of the effects results from the environment-induced thermalization of the qubit system in the course of the driving procedure, which under some circumstances increases population of the target ground state. Another effect follows from an initial thermalization of the system at the start of the driving procedure, which for a certain optimal tempertature leads to an efficient transfer of populations from excited states to the ground state in the QPT and ESQPT regions. Both effects are present for drivings across the first- as well as second-order QPT. The latter effect was nevertheless found to weaken with an increasing size of the system, which limits its application in driving procedures involving large numbers of qubits.
@@ ACKNOWLEDGEMENTS
We thank Jakub Dolejší for his help in initial stages of this work within his diploma thesis Dole20 .

### v4

@@ 4.3 Comparison with the Lindblad method
Figure 9: Fidelity for driving A across the first-order QPT (with ${\theta=0}$ and ${r=0}$) for the ${N=10}$ qubit system obtained by the HEOM method (full curves) and by the Lindblad method according to Ref. Yama17 (dashed curves). The driving times $t_{\rm F}$ and coupling strengths $q$ are indicated, the range of temperature is $\log_{10}T\in[-0.8,1.4]$.
@@
In the comparison of the Lindblad calculations with the HEOM results, we consider only the path across the first-order QPT with the driving protocol A (see Sec. 4.1). The S-E coupling operator (8) has ${\theta=0}$ and the counterterm in (9) is not included (${r=0}$). If the dynamics for ${N=10}$ qubits is modeled using the two-state approximation, the adiabatic timescale is of the order ${t=2|d\Delta_{10}/ds|/\pi\Delta_{10}^{2}\sim 10^{3}}$, where all terms are evaluated at the crossing Dick13 . The method of Ref. Alba12 ; Yama17 is limited to nearly-adiabatic drivings, so only the drivings with ${t_{\rm F}\gg 10^{3}}$ can be expected to agree with the Lindblad calculations.
In Fig. 9 we show the fidelity $\mathcal{F}_{1}$ from Eq. (25) obtained by both the HEOM and Lindblad methods for driving times $t_{\rm F}=10^{2.8}$ and $10^{3.6}$. We see, in accord with the above expectation, that for the larger $t_{\rm F}$ the discrepancy between both methods affects only the drivings at lower temperatures ($T\lesssim\gamma/2\pi$, see Ref. Tani20 ) and shows up only for the strong S-E coupling. For this driving time, the Lindblad approach predicts the final state of the system that essentially coincides with the canonical equilibrium state $\hat{\rho}^{(\beta)}_{\rm S}(\mathbb{\Lambda}_{\rm F})\propto e^{-\beta\hat{H}_{\rm S}(\Lambda_{\rm F})}$. The discord of the HEOM calculations with this prediction for the stronger coupling and low temperatures agrees with the results of Refs. Dijk12 ; Iles14 ; Iles16 , where deviations of the actual final state from the canonical state were also observed. Note that a realistic approximation for the equilibrium state for all coupling strengths was derived in Refs. Anto23_1 ; Anto23_2 . For the shorter $t_{\rm F}$ in Fig. 9, the HEOM and Lindblad results differ for both coupling strengths and all temperatures below the high-$T$ limit (28).
@@
Therefore, we can conclude that non-Markovian effects, properly captured in our HEOM calculations, play in general an important role, which is not reproduced by the simpler Lindblad calculations.
@@ 5 Conclusion
However, thermalization and dissipation processes can strongly modify results of the driving procedure, in some cases being able to considerably improve its fidelity. One of the effects results from the environment-induced thermalization of the qubit system in the course of the driving procedure, which under some circumstances increases population of the target ground state. Another effect follows from an initial thermalization of the system at the start of the driving procedure, which for a certain optimal temperature leads to an efficient transfer of populations from excited states to the ground state in the QPT and ESQPT regions. Both effects are present for drivings across the first- as well as second-order QPT. The latter effect was nevertheless found to weaken with an increasing size of the system, which limits its application in driving procedures involving large numbers of qubits.
@@ ACKNOWLEDGEMENTS
We thank Jakub Dolejší for his help in the initial stages of this work within his diploma thesis Dole20 .